Operations Report – January 10th

Operations Report

Crew 219 Operations Report 10-JAN-2020

Sol: 5

Name of person filing report: Alejandro Perez

Non-nominal systems: N/A

Generator: Run

Hours run: 14

From what time last night: 17:00

To what time this morning: 07:00

List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before the generator is run at night): 84%

10:30 =100%, 11:00 =99%, 11:30 = 98%, 12:00 =97%, 12:30= 96%, 13:00 =
95%, 13:30 = 93%, 14:00 = 91%, 14:30 = 90%, 15:00 = 89%, 15:30 = 88%,
16:00 = 87%, 16:30 = 87%, 17:00 = 84%

Diesel Reading – 70%

Station Propane Reading- 45%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 50

Water Meter (units)- 0147538,3

Water (static tank) (gal) – 468

Static to loft Pump used – Yes

Water in Green Hab (gal): 188.65

Water in Science Dome (gal): 0

Toilet Tank Emptied – No

Diemos rover used: Still in the Work shop

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: Assigned to Director

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used- No

Starting Hours:

Beginning Charge:

Ending Hours:

Ending Charge:

Opportunity Rover used: Still in workshop

Hours:

Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: No

Starting Hours:

Beginning Charge:

Ending Hours:

Ending Charge:

Notes on Rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? No

# hours the ATV’s were used today:

Notes on ATVs:

Habcar used and why, where? Used by Outpost Team to go to the storage
room, get water and
supplies for Outpost and Hab.

CrewCar used and why, where? Off-Campus, still on the workshop for repair.

General Notes and Comments:

Summary of the internet: Nominal

Summary of Suit and Radios: Defective radios have been placed inside
the RAM for Outpost evaluation. The repairs for suit #3 was tested on
EVA 10 today and performed perfectly. All two-piece helmets are
missing one O-Ring for their connectors.

Summary of Hab: Pipe froze again today. The heating coil is still
functioning and the pump is still operating properly. The freezing
seems to be in between the exterior wall and the interior wall leading
to the washroom. This was solved using a combined method of warm water
through the feedline, a hair dryer heating the interior. The
thermostat in the was room read lowest temperature at -2.0°C during
the night. Outpost did a repair today to the exterior section of the
pipe. On the 9th, solar panels were inspected for any damage. All the
system and structure are nominal and fully operational. Heat tape for
the pipelines from the static tank to the Hab is being rearranged in
order to completely fix the frozen pipes issue. Furnace filter was
also changed today. We received Teflon tape, toilet paper, and pipe
insulation.

Summary of Science Dome operations: Temperatures are hovering around 7
°C during the day.

Summary of Ram operations: Nominal

Summary of any observatory issues: Nominal

Summary of Health and safety issues: Nominal

Questions, concerns, and requests to Mission Support: 11 O-Rings (7
replacements and 4 Bench stock) and new transformer for radio charging
unit (input=120VAC 60Hz 9W, Output=9V DC 300 mA)

Journalist Report – January 10th

Crew 219 Journalist Report 10 JAN 2020

Nathan Hadland, Crew 219 Executive Officer

Sol 5

A Tangy Day

The morning started out like any other; people started climbing out of their beds and making coffee and tea. However, the environment outside was anything but normal. A thick layer of icy fog stretched over the landscape. On the tunnels around the Hab, there was a coat of freezing fog which was really beautiful in the sunlight. Until about 1000 hours, visibility was extremely limited, so we decided to scrap our morning EVA. However, once the sun was higher in the sky, the fog disappeared in a span of 15 minutes! It was as if a vacuum had rolled over the land and sucked it all up! We were left with blue skies for the first time our entire mission. These clear skies have us hopeful that we will finally get some data for our astronomy research.

The scrapped EVA freed up our morning to catch up on work in the lab and some maintenance issues. I helped Alejandro (ENG) thaw the frozen pipes and fix some of the EVA suits. Meanwhile, Cynthia (GHO) and Abdul (GEO) worked in the lab on their samples. Unfortunately, it looks like our cyanobacteria have died due to the extreme temperature fluctuations they have been subjected to. However, we will continue to analyze our samples and their effect on the morphology and mineralogy of the regolith.

Due to the incredible change in weather in the morning, the second EVA was good to go. I led Alejandro (ENG), Abdul (GEO), and Robinson (ASTRO) up Hab Ridge and the views were breathtaking. We could clearly see Mt. Pennell, Skyline Rim, among other amazing geologic features. We made the long trek through the snow along Sagan Road to Mid Ridge Planitia to get samples for our mineralogy survey. Tomorrow, we will continue this survey by going all the way to Skyline Rim to take samples at the base of the mountain. Along the way, we will be keeping our eyes out for lichen or other microbiological samples for our astrobiology project.

Although previous journalist reports have commented on the excellency of Hannah’s (LSO) cooking, I feel it is necessary to continue these praises. Last night she made Tang-flavored teriyaki chicken and today she made Tang-flavored cinnamon rolls. Each bite I took, I felt like the universe was speaking to me through food and unlocking its mysteries.

Research Report – January 10th

Crew 219 Science Report 10-JAN-20
Crew Science Officer: Hannah Blackburn

1. Biometrics and Neurobehavioral Research

Continued taking sleep and post-EVA logs. Post-EVA biometrics were recorded.

2. UAV

Nothing to report.

3. Dust Mitigation for Optical Mirrors

Mirrors were recovered during EVA 10 and returned to the Hab for analysis.

4. Astrophotography of Celestial Bodies

Nothing to report. Still poor visibility 🙁

5. Remediation of Mars Regolith

Inoculated regolith examined under microscope. Cyanobacteria appears
to have died, most likely due to wide fluctuations in temperature.

6. Chemical and Mineralogical Composition of the MDRS Site

Samples were dried in the oven using glass beakers. Additional samples
were taken during EVA 10.

7. Protocols for the Discovery of Life on Mars

Nothing to report.

Glassware check out:

10 glass beakers (sizes between 200 and 500 ml) were used to dry
regolith samples.

GreenHab Report – January 10th

Crew 219 GreenHab Report 10-JAN-20

Crew GreenHab Officer: Cynthia Montanez

Environmental control: Heating.

Shade cloth (40% and 30%) on.

Average temperature: 24 °C; 18 %

Average Max: 29.3 °C, 21%

Average Min: 17 °C, 15.5%

9:40 A.M.

Floor Unit: 4 °C

Electronic: 18.8 °C

humidity 20%

Max: 27.4 °C; 24%

Min: 11.7 °C; 16%

12:58 P.M.

Floor Unit: 20 °C

Electronic: 28.4 °C

humidity 16%

Max: 28.7 °C; 20%

Min: 18.8 °C; 16%

2:56 P.M.

Floor Unit: 14 °C

Electronic: 24.7 °C

humidity 18%

Max: 30.3 °C; 20%

Min: 18.8 °C; 15%

4:52 P.M.

Floor Unit: 14 °C

Electronic: 24.2 °C

humidity 18%

Max: 30.7 °C; 20%

Min: 18.8 °C; 15%

Hours of supplemental light: Light system 7:00PM-12:00AM.

Daily water usage for crops: 4.46 Gal.

Daily water usage for research and/or other purposes: N/A

Water in Blue Tank – ~188.65 Gal.

Time(s) of watering for crops:

9:40 A.M.

SEASONAL PLANTING HISTORY:

Change to crops: The icicle radish has a lot of bite marks on the leaves and small mites flying around it when the leaves are touched. I am assuming that there is an infestation of some bug, e.g. caterpillar, eating the plant. I sprayed the plant with organic pesticide to attempt to kill the bugs.

OTHER NOTES:

1st sprouts: N/A

Harvest: N/A

Other: Since the thermostat is set to 21 , the GreenHab has gotten substantially hotter. I needed to leave the door open many times so that the temperature would not exceed 30 . This increase in temperature is still negatively affecting the plants by the GreenHab door by making the leaves crispy and the soil dry out faster.

Suggestions: After a brief discussion with Outpost, I was wondering if it was possible to move the thermostat to a cooler 19 . The temperature would not be too cold for the plants, and it would not create too much heat resulting in the door needing to be opened constantly.

Sol Summary – January 10th

Crew 219 Sol Summary Report 10-01-2020

Sol: 5

Summary Title: Frozen Fog & Tang-glazed Cinnamon Rolls

Author’s name: Dave Masaitis

Mission Status: Nominal

Sol Activity Summary: The morning broke with the entire valley
shrouded in thick fog, dropping visibility next to nothing. By the
time EVA #9 was intended to depart, the morning sunshine had burned
off much of the fog, but ground conditions did not appear to support
safe maneuver conditions, so we decided to cancel it. The crew instead
busied themselves with science and maintenance, while H. Blackburn
prepared a wonderful lunch of pork sausage chili and fresh-baked
cornbread. The team for EVA #10 ate lunch and moved out on schedule,
achieving both their maneuver plan and sampling schedule. On the way
back, they recovered R. Raphael’s experimental mirror emplacement for
analysis. Upon their return they found another surprise, courtesy of
H. Blackburn – steaming hot chocolate and Tang-glazed cinnamon rolls –
a real treat! With full bellies, some crew members took a nap, while
others refined their reports for the evening Comms window. This
evening will consist of relaxed socialization while crew members
recoup for another Sol of hard work tomorrow.

Look Ahead Plan: We anticipate two EVA’s, and will continue ongoing
scientific analysis and maintenance operations.

Anomalies in work: None

Weather: Cold and foggy in the early morning, followed by clear,
sunny, and frozen conditions the rest of the day

Crew Physical Status: Nominal

EVA:

EVA #9 – Cancelled at Commander’s Discretion

EVA #8 – Geologic Sample Collection South of Mid Ridge Planitia
(0516800/4251600)

Reports to be filed:

Sol Summary

EVA Report 10

EVA Requests 11 & 12

Operations Report

GreenHab Report

Research Report

Journalist’s Report

Photos (Including Photo of the Day)

Support Requested: Refer to Operations Report

EVA Report – January 10th

Crew 219 EVA Report 10 JAN 2020

EVA #10

Author: Nathan Hadland

Purpose of EVA: The primary purpose of this EVA was sampling for our
geology project. Post-EVA, we collected data for our biometrics
project.

Start time: 1253 hrs

End time: 1513 hrs

Narrative: EVA #10 initiated a 5-minute decompression at 1253 hours
and headed north towards North Ridge. We then proceeded up Hab Ridge
using the Sagan Street (1103) access road. Abdul Elnagdi (GEO)
identified a site for sampling and we took 5 samples at the top of the
ridge and recorded GPS locations. The EVA team then walked along Sagan
Street until reaching Mid Ridge Planitia and took 5 additional
samples. The team then proceeded back along Sagan Street and descended
Hab Ridge. The team arrived safely back to the Hab and picked up the
optical mount for analysis. The EVA crew entered the air lock at 1508
hours and successfully pressurized in the air lock at 1513 hours.

Destination: Just east of Mid Ridge Planitia

Coordinates: 0516800/4251600

Participants: Nathan Hadland (XO), Alejandro Perez (ENG), Robinson
Raphael (ASTRO), Abdul Elnagdi (GEO)

Roads and routes per MDRS Map: North towards North Ridge and then
climbed Hab Ridge near Schubert Pass using Sagan Street (1103). We
then headed west towards Mid Ridge Planitia following Sagan Street
(1103).

Mode of travel: Walking